Graph the ellipse x^2+4y^2=64
Please help. Thank you.

Answer by lwsshak3(11628)

Graph the ellipse
x^2+4y^2=64
4y^2=(64-x^2)
y^2=(64-x^2)/4
y=±(64-x^2)^.5/2
see graph below: